jill is 0
bill is 28
how old will they be when bill is 3 times her age?

Answer by Theo(13342)

let x = jill's age.
then x + 28 = bill's age.
let y = the number of years it will take for bill's age to be equal to 3 * jill's age.

you get (x + 28) + y = 3 * (x + y)

since x = 0, the equation becomes:

28 + y = 3 * y

subtract y from both sides of this equation to get:

28 = 2 * y

divide both sides of this equation by 2 to get:

y = 14

in 14 years, bill's age will be 3 * jill's age.

bill will be 28 + 14 = 42
jill will be 14

3 * 14 = 42

solution is:

bill will be 42
jill will be 14.